OpenStudy (anonymous):

Evaluate the following limit . lim x->5- lnx/(x-5)^2

OpenStudy (lgbasallote):

are you familiar with the l'hospital rule?

OpenStudy (anonymous):

nah

OpenStudy (anonymous):

you cannot use l'hopital here

OpenStudy (anonymous):

the denominator goes to zero, but the numerator does not, so no limit

OpenStudy (anonymous):

I believe it's infinity. \[\begin{align}\lim_{x\rightarrow 5} \ln x = \ln 5\\ \lim_{x\rightarrow 5} (x-5) = 0\end{align}\] Hence you have a positive real number over a very, very small positive real (the square ensures the denominator is positive), implying the limit is infinity.